A set of data with a mean of 58 and a standard deviation of 6.3 is normally distributed. Find the value that is 2 standard deviations below the mean. Round your answer to the nearest tenth.
Answer by jim_thompson5910(35256)

Simply subtract 2 times the standard deviation from the mean to get:


58-2*6.3=58-12.6=45.4


So the value that is 2 standard deviations below the mean is 45.4